# DiffusionNFT: Online Diffusion Reinforcement with Forward Process

### v1

@@ Appendix C Experiment Details
Multi-Reward. To comprehensively improve the base model across multiple rewards, we adopt a multi-stage training scheme. The training setup involves three categories of rewards and datasets: (1) PickScore, CLIPScore, and HPSv2.1 on the Pick-a-Pic dataset; (2) GenEval with the three rewards above on the GenEval dataset; and (3) OCR with the three rewards above on the OCR dataset. Since the initial CFG-free generation is of low quality, we first train on (1) for 800 iterations to enhance image quality, followed by (2) for 300 iterations, (1) for 200 iterations, (2) for 200 iterations, and finally (3) for 100 iterations. All rewards are equally weighted, with PickScore divided by 26 for normalization to $[0,1]$. By default, we use $\beta=0.1$ and $\eta_{i}=\min(0.001i,0.5)$, while setting $\eta_{\max}=0.95$ for OCR to stabilize training. The number of sampling steps is fixed to 40 to ensure high-fidelity data collection.

### v2

@@ Appendix C Experiment Details
Multi-Reward. To comprehensively improve the base model across multiple rewards, we adopt a multi-stage training scheme. The training setup involves three categories of rewards and datasets: (1) PickScore, CLIPScore, and HPSv2.1 rewards on the Pick-a-Pic dataset; (2) GenEval reward with the three rewards above on the GenEval dataset; and (3) OCR reward with the three rewards above on the OCR dataset. Since the initial CFG-free generation is of low quality, we first train on (1) for 800 iterations to enhance image quality, followed by (2) for 300 iterations, (1) for 200 iterations, (2) for 200 iterations, and finally (3) for 100 iterations. All rewards are equally weighted, with PickScore divided by 26 for normalization to $[0,1]$. By default, we use $\beta=0.1$ and $\eta_{i}=\min(0.001i,0.5)$, while setting $\eta_{\max}=0.95$ for OCR to stabilize training. The number of sampling steps is fixed to 40 to ensure high-fidelity data collection.
@@ Appendix D Additional Results
Table 2: Evaluation results of FlowGRPO and DiffusionNFT trained on single rewards, both initialized from CFG-free base model. Gray-colored: In-domain reward. We observe that training exclusively on the OCR reward impairs generalization to other metrics; to compensate this, we enable CFG when evaluating non-OCR rewards for OCR-trained models.
Model #Iter Rule-Based Model-Based
GenEval OCR PickScore ClipScore HPSv2.1 Aesthetic ImgRwd UniRwd
SD3.5-M (w/o CFG) 0.24 0.12 20.51 0.237 0.204 5.13 -0.58 2.02
+ CFG 0.63 0.59 22.34 0.285 0.279 5.36 0.85 3.03
+ FlowGRPO 4k 0.97 0.30 21.78 0.277 0.248 5.15 0.74 2.87
1k 0.66 0.96 21.94 0.280 0.257 5.18 0.31 2.86
4k 0.54 0.60 23.62 0.257 0.295 6.42 1.17 3.17
+ Ours 1k 0.98 0.36 21.92 0.271 0.251 5.33 0.68 2.91
150 0.54 0.97 21.63 0.281 0.246 5.19 0.37 2.81
2k 0.53 0.64 24.03 0.270 0.315 6.17 1.29 3.40